2(5a-2)+3(4-7a)=-25
Answer by BGluvsmath(64)

2(5a-2)+3(4-7a)=-25
Distribute: 10a-4+12-21a=-25
Simplify: -11a+8=-25
subtract 8 from both sides: -11a+8-8=-25-8, simplified to -11a=-33
divide each side by -11: -11a/-11 = -33/-11, simplified to a=3